OpenStudy (credmond):

Let z = sqrt{3x+2y}. Then: The rate of change in z at (1,1) as we change x but hold y fixed is___ and The rate of change in z at (1,1) as we change y but hold x fixed is___

OpenStudy (jhannybean):

\[z = \sqrt{3x+2y}\]\[\frac{\partial x}{\partial z} = ~?\]\[\frac{\partial y}{\partial z} = ~?\]

OpenStudy (jhannybean):

ell me what those two are first. Then we'll continue :D

OpenStudy (credmond):

So x with respect to z and y with respect to z, right?

OpenStudy (credmond):

partial derivative with respect...

OpenStudy (jhannybean):

Yep, because z is your function.

OpenStudy (jhannybean):

if we're changing x, then x is our variable and y is our constant. if we're changing y, then y is our variable and x is our constant

OpenStudy (credmond):

Okay!

OpenStudy (jhannybean):

Just a min.

OpenStudy (jhannybean):

\[\frac{\partial x}{\partial z} = \frac{1}{2}(3x+2y)^{-1/2}(3) = \frac{3}{2(3x+2y)^{1/2}}\]

OpenStudy (jhannybean):

So now when you input (1,1) for (x,y), what do you get as your value for \(\dfrac{\partial x}{\partial z}\)?

OpenStudy (anonymous):

it think it should be 

OpenStudy (credmond):

Oh okay! So you get 3/2 sqrt 5?

OpenStudy (jhannybean):

@surjithayer you are right

OpenStudy (jhannybean):

Yes that is what i got @credmond

OpenStudy (jhannybean):

you do the next one

OpenStudy (credmond):

Sooo... Is it just 2/ 2 (3x+2y)^1/2 ?

OpenStudy (credmond):

which gives you 2/2 sqrt (5)

OpenStudy (jhannybean):

There you go \(\checkmark\)

OpenStudy (credmond):

Aaaaaah! Thank you sooooo much I really appreciate the help! You are the best! :)